### v3

# Graph quandles: Generalized Cayley graphs of racks and right quasigroups

Abstract.
We solve two open problems of Valeriy Bardakov about Cayley graphs of racks and graph-theoretic realizations of right quasigroups. We also extend Didier Caucal’s classification of labeled Cayley digraphs to right quasigroups and related algebraic structures like quandles.
First, we characterize markings of graphs that realize racks. As an application, we construct rack-theoretic (di)graph invariants from permutation representations of graph automorphism groups. We describe how to compute these invariants with general results for path graphs and cycle graphs.
Second, we show that all right quasigroups are realizable by edgeless graphs and complete (di)graphs. Using Schreier (di)graphs, we also characterize Cayley (di)graphs of right quasigroups $Q$ that realize $Q$. In particular, all racks are realizable by their full Cayley (di)graphs.
Finally, we give a graph-theoretic characterization of labeled Cayley digraphs of right-cancellative magmas, right-divisible magmas, right quasigroups, racks, quandles, involutory racks, and kei.
Key words and phrases:
Cayley graph, groupoid, labeled directed graph, magma, quandle, rack, right quasigroup
2020 Mathematics Subject Classification:
Primary 05C25, 20N05, 57K12; Secondary 05C75, 05C78, 20N02
1. Introduction
Right quasigroups vastly generalize racks and quandles, which are nonassociative algebraic structures used to construct invariants in group theory, knot theory, and low-dimensional topology. During a 2020 lecture, Bardakov introduced a way to construct right quasigroups from markings of graphs by graph automorphisms [bardakov]. Interested in finding geometric interpretations of racks and quandles, Bardakov closed the lecture by posing two open questions about which marked graphs realize racks and how they relate to Cayley graphs of racks; see Problems 1.6 and 1.7.
In this paper, we answer Bardakov’s questions with the following results. We also introduce two rack-theoretic (di)graph invariants $\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}}$ and characterize labeled Cayley digraphs of racks and quandles. We refer to directed graphs as digraphs and simple undirected graphs as graphs.
Theorem 1.1.
Let $\Gamma$ be a (di)graph with vertex set $V$, and let $R:V\to\operatorname{Aut}\Gamma$ be a marking (resp. q-marking) of $V$. Then the right quasigroup $V^{\Gamma}_{R}$ realized by $\Gamma$ is a rack (resp. quandle) if and only if $R$ is a magma homomorphism from $V^{\Gamma}_{R}$ to $\operatorname{Conj}(\operatorname{Aut}\Gamma)$. (See Theorem 5.1.)
Proposition 1.2.
All right quasigroups are realizable by edgeless graphs and complete (di)graphs. (See Proposition 3.9.)
Theorem 1.3.
Let $V_{R}$ be a right quasigroup, and let $\Gamma$ be a Cayley digraph (resp. graph) of $V_{R}$ with connection set $S\subseteq V$. Then $(\Gamma,R)$ realizes $V_{R}$ if and only if, for all $h,v\in V$ and $s\in S$, there exists an element $t\in S$ such that
$R_{t}R_{h}(v)=R_{h}R_{s}(v)$
(resp. $R_{t}^{\pm 1}R_{v}(w)=R_{v}R_{s}(w)$). (See Theorems 6.16.2.)
Corollary 1.4.
All racks are realizable by their full Cayley (di)graphs. (See Corollary 6.4.)
To our knowledge at the time of writing, this paper is the first to study Bardakov’s constructions. It is also the first to characterize labeled generalized Cayley digraphs of right quasigroups, racks, and quandles, as summarized below; cf. Problem 1.8. This extends Caucal’s work on labeled Cayley digraphs of quasigroups, left quasigroups [caucal], groups, monoids, and semigroups [caucal2].
Theorem 1.5.
Let $\mathcal{Q}$ be the class of all labeled digraphs that are deterministic, source-complete, codeterministic, and target-complete. Then $\mathcal{Q}$ is precisely the class of labeled Cayley digraphs of right quasigroups. (See Theorem 7.12.)
Moreover, there exist graph theoretic-conditions on $\mathcal{Q}$ that restrict it to the subclass of labeled Cayley digraphs of racks, and similarly for quandles, involutory racks, and kei. (See Theorem 7.14 and Corollary 7.15.)
1.1. Motivating discussion
We give an overview of the historical background and open questions motivating this work. See Section 2 for formal definitions of the terms discussed here.
1.1.1. Quandle theory and generalizations
[…]
Generalizations of quandles have also been of interest in many areas of mathematics. In particular, Fenn and Rourke [fenn] introduced racks in 1992 to construct complete invariants of closed, connected 3-manifolds and framed links embedded in them. Racks are also used to study Hopf algebras (e.g., [hopf]), the Yang–Baxter equation (see [yang]), and Legendrian knots in contact three-space (see [ta]), for instance. On the other hand, right quasigroups are a class of magmas or groupoids that vastly generalizes racks, quandles, groups, and quasigroups. Various authors have used right quasigroups to study column Latin squares [Latin], smooth deformations of Lie group structures [diff], and nonassociative generalizations of Hopf algebras [quantum], for instance.
1.1.2. Racks as symmetries
Because of the myriad uses of racks and quandles, novel ways of understanding their structure are desirable. Inspired by a class of quandles called dihedral quandles, which model reflections of cycle graphs, Bardakov sought geometric interpretations of other classes of racks and quandles during a 2020 lecture [bardakov]. To formalize this search, he introduced realizations of right quasigroups by marked graphs, which are graphs $\Gamma$ with an assignment $R$ of each vertex to a graph automorphism. Bardakov closed his lecture with the following open problems.
Problem 1.6.
Under what conditions does a marked graph realize a rack or a quandle?
Problem 1.7.
Given a rack or quandle $Q$, is there always a marked graph $(\Gamma,R)$ that realizes $Q$? If so, can we choose $\Gamma$ to be a Cayley graph of $Q$?
Theorem 5.1 answers Problem 1.6. Since all racks are right quasigroups, Proposition 3.9 answers a generalized version of the first question in Problem 1.7. Corollary 6.4 answers the second question in Problem 1.7, while Theorems 6.16.2 answer generalized forms of the question.
1.1.3. Characterization of labeled Cayley digraphs
Although this paper is the first to study labeled Cayley (di)graphs of right quasigroups in general, various authors have studied Cayley graphs of various classes of right quasigroups appearing in combinatorics, algebraic topology, and knot theory. For example, full Cayley digraphs of unital, fixed point-free right quasigroups have various applications in network theory [groupoid1], and full Cayley graphs of right quasigroups have applications in categorical covering theory [categories].
[…]
Problem 1.8.
Given a full subcategory $\mathcal{C}$ of magmas (e.g., groups, monoids, quandles), are there graph-theoretic conditions that characterize labeled Cayley digraphs of objects in $\mathcal{C}$?
Caucal answered Problem 1.8 for left quasigroups, quasigroups [caucal], semigroups, and various classes of monoids, including groups [caucal2]. [^2] Since the right-multiplication maps $R_{v}$ of left quasigroups are not necessarily permutations, the answers to Problem 1.8 for left and right quasigroups are distinct. Chishwashwa et al. addressed a similar question for vertex-labeled Cayley digraphs of unital, fixed point-free right quasigroups [groupoid1]. In this paper, we answer Problem 1.8 for various classes of racks and right quasigroups; see Theorems 7.12 and 7.14 and Corollary 7.15.
1.2. Structure of the paper
In Section 2, we discuss right quasigroups, racks, and quandles.
In Section 3, we discuss (di)graphs, marked graphs in the sense of Bardakov, Cayley (di)graphs of magmas in the sense of Caucal, and Schreier graphs of group actions.
In Section 4, we give examples of Cayley (di)graphs of right quasigroups and their markings.
In Section 5, we answer Problem 1.6. As an application, we introduce two rack-theoretic (di)graph invariants $\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}}$ with general results for path graphs and cycle graphs.
In Section 6, we answer Problem 1.7 and its analogues for right quasigroups and digraphs.
In Section 7, we define labeled Cayley digraphs of magmas in the sense of Caucal, and we answer Problem 1.8 for right-cancellative magmas, right-divisible magmas, right quasigroups, racks, quandles, involutory racks, and kei.
In Section 8, we propose directions for future research.
1.3. Notation
For all positive integers $n\in\mathbb{Z}^{+}$, let $[n]$ denote the set $\{1,2,\dots,n\}$. Denote the symmetric group of $[n]$ by $S_{n}$ with its elements written in cycle notation, and denote the symmetric group of any other set $X$ by $S_{X}$. We also denote the composition of functions $\varphi:V\to W$ and $\psi:W\to X$ by $\psi\varphi$, and we denote the identity map on a set $V$ by $\operatorname{id}_{V}$. For $n\geq 3$, let $D_{n}$ be the dihedral group of order $2n$. Given a subset $S$ of a group $G$, let $S^{-1}:=\{s^{-1}\mid s\in S\}$.
[…]
$v\triangleright w=R_{w}(v).$
2. Algebraic preliminaries
We recall the definitions of right quasigroups, racks, and quandles.
2.1. Magmas
Right quasigroups are examples of more general algebraic structures called magmas.
Definition 2.1.
A magma or groupoid is a pair $(V,R)$, denoted by $V_{R}$, where $V$ is a set and $R$ is a mapping from $V$ to the set of functions from $V$ to VV. [^3] Although the notation $V_{R}$ is nonstandard, we use it because Bardakov formulated his open problems with it [bardakov]. For all $v\in V$, we call the map $R_{v}:=R(v)$ a right-multiplication map or right-translation map.
Moreover, we say that $V_{R}$ is a right quasigroup if $R(V)\subseteq S_{V}$, that is, if all right-multiplication maps are permutations of $V$. We say that $R$ is a right quasigroup structure on $V$.
Example 2.2.
The (right) regular action $R:G\to S_{G}$ of a group $G$, given by $R_{h}(g):=gh$, is a right quasigroup structure on $G$. Thus, right quasigroups generalize groups.
Example 2.3.
If we take $G:=\mathbb{Q}$ to be the rational numbers in Example 2.2, then the positive rational numbers $\mathbb{Q}^{+}$ are a submagma of $\mathbb{Q}$ but not a right quasigroup.
Definition 2.4.
Let $V_{R}$ and $W_{T}$ be magmas. A magma homomorphism from $V_{R}$ to $W_{T}$ is a function $\varphi:V\to W$ satisfying
$\varphi R_{v}=T_{\varphi(v)}\varphi$
for all $v\in V$.
With this definition, magmas and right quasigroups form categories. In particular, we can consider automorphism groups $\operatorname{Aut}V_{R}$ of magmas.
2.2. Racks
Racks and quandles form important full subcategories of the category of right quasigroups.
Definition 2.5.
Let $V_{R}$ be a right quasigroup.
  • The right-multiplication group $\operatorname{RMlt}V_{R}$ is the subgroup of $S_{V}$ generated by all right-multiplication maps.
  • We say that $V_{R}$ is a rack if every right-multiplication map $R_{v}$ is a magma endomorphism. Concretely, this means that
    (2.1) $R_{v}R_{w}=R_{R_{v}(w)}R_{v}$
    for all $v,w\in V$. In this case, we call $R$ a rack structure on $V$.
  • Separately, we say that $V_{R}$ is involutory if every right-multiplication map is an involution, that is, if $R_{v}^{2}=\operatorname{id}_{V}$ for all $v\in V$.
Remark 2.6.
A right quasigroup $V_{R}$ is a rack if and only if $\operatorname{RMlt}V_{R}$ is a (normal) subgroup of $\operatorname{Aut}V_{R}$. In this case, some authors call $\operatorname{RMlt}V_{R}$ the inner automorphism group of $V_{R}$ and denote it by $\operatorname{Inn}V_{R}$. Other authors denote $\operatorname{RMlt}V_{R}$ by $\operatorname{Mlt}_{r}V_{R}$.
Definition 2.7.
Let $V_{R}$ be a rack.
  • We say that $V_{R}$ is a quandle if $R_{v}(v)=v$ for all $v\in V$. In this case, we call $R$ a quandle structure on $V$.
  • If $V_{R}$ is an involutory quandle, we call it a kei.
2.2.1. Examples
We discuss some common examples of right quasigroups, racks, and quandles. See Section 4 for further examples.
Example 2.8.
[book]*Example 2.13 Let $G$ be a union of conjugacy classes in a group, and define $C:G\rightarrow S_{G}$ by sending any element $g\in G$ to the conjugation map $C_{g}$ defined by
$C_{g}(h):=ghg^{-1}.$
Then $\operatorname{Conj}G:=G_{C}$ is a quandle called a conjugation quandle or conjugacy quandle.
If $G$ is a group, then $\operatorname{Conj}G$ is a kei if and only if $g^{2}\in Z(G)$ for all $g\in G$. In particular, not all quandles are involutory.
Example 2.9.
[quandlebook]*Example 99 Let $V$ be a set, fix a permutation $\sigma\in S_{V}$, and define $R_{v}:=\sigma$ for all $v\in V$. Then the assignment $R$ is a rack structure on $V$, and we call $V_{\sigma}:=V_{R}$ a permutation rack or constant action rack.
Note that $V_{\sigma}$ is a quandle if and only if $\sigma=\operatorname{id}_{V}$, in which case we call $V_{\operatorname{id}_{V}}$ a trivial quandle. In particular, not all racks are quandles. Moreover, $V_{\sigma}$ is involutory if and only if $\sigma$ is an involution.
Example 2.10.
The (right) regular action of a group $G$ is a rack structure if and only if $G$ is the trivial group. In particular, not all right quasigroups are racks.
2.2.2. Preliminary results
An alternative characterization of racks does the heavy lifting in solving Problem 1.6; see Theorem 5.1.
Proposition 2.11.
Let $V_{R}$ be a magma. Then $V_{R}$ is a rack if and only if $R$ is a magma homomorphism from $V_{R}$ to $\operatorname{Conj}S_{V}$.
Proof.
$\implies$” Suppose that $V_{R}$ is a rack. Then $R(V)\subseteq S_{V}$, and Equation (2.1) states that
$RR_{v}(w)=R_{R_{v}(w)}=R_{v}R_{w}R_{v}^{-1}=C_{R_{v}}(R_{w})=C_{R(v)}R(w)$
for all $v,w\in V$. Hence, $R$ is a magma homomorphism from $V_{R}$ to $\operatorname{Conj}S_{V}$.
[…]
$R_{R_{v}(w)}=RR_{v}(w)=C_{R(v)}R(w)=C_{R_{v}}(R_{w})=R_{v}R_{w}R_{v}^{-1}$
for all $v,w\in V$. Since $R_{v}\in S_{V}$ is bijective, we obtain Equation (2.1). ∎
In Sections 4 and 6, we employ a necessary condition for a right quasigroup to be a rack.
Lemma 2.12.
If $V_{R}$ is a rack, then $R(V)$ is closed under conjugation.
Proof.
Equation (2.1) states that
$R_{v}R_{w}R_{v}^{-1}=R_{R_{v}(w)}\in R(V)$
for all $v,w\in V$. ∎
3. Graph-theoretic preliminaries
We discuss directed and undirected graphs, marked graphs as constructed by Bardakov [bardakov] and Cayley graphs of magmas as introduced by Caucal [caucal]. We also relate the latter to Schreier graphs of group actions. (Since we only discuss labeled digraphs in Section 7, we defer defining them until then.)
3.1. Graphs and digraphs
We recall the graph-theoretic constructions of Bardakov in [bardakov]. Like Bardakov, we assume that all undirected graphs are simple, and we do not allow for digraphs to have multiple edges. However, we do allow for digraphs to have loops.
Definition 3.1.
  • A digraph or directed multigraph $\Gamma$ is a pair $(V,E)$ where $V$ is a set and $E$ is a subset of $V\times V$. We say that $V$ is the vertex set of $\Gamma$, and we say that $E$ is the (directed) edge set or arc set of $\Gamma$, respectively.
  • Simple undirected graphs, which we only call graphs, are defined similarly to digraphs, except that every element of $E$ is an unordered pair of vertices $\{v,w\}\subseteq V$ such that $v\neq w$.
  • Given a (di)graph $\Gamma$, we denote the vertex and edge sets of $\Gamma$ by $V(\Gamma)$ and $E(\Gamma)$, respectively. We say that $|V(\Gamma)|$ is the order of $\Gamma$.
Definition 3.2.
Let $\Gamma=(V,E)$ be a digraph. An automorphism of $\Gamma$ is a permutation $\varphi\in S_{V}$ such that $(\varphi(v),\varphi(w))\in E$ for all edges $(v,w)\in E$. Automorphisms of graphs are defined similarly.
As with racks and quandles, digraphs and graphs form categories. In particular, we can consider automorphism groups $\operatorname{Aut}\Gamma$ of digraphs and graphs.
3.2. Marked graphs
We consider (di)graphs with markings and q-markings of their vertices as introduced by Bardakov [bardakov].
Definition 3.3.
Let $\Gamma$ be a (di)graph with vertex set $V$.
  • A marking of $\Gamma$ is a function $R:V\to\operatorname{Aut}\Gamma$ with each image notated as $R_{v}:=R(v)$. We say that the right quasigroup $V_{R}^{\Gamma}:=(V,R)$ is realized by the marked (di)graph $(\Gamma,R)$.
  • Let $R$ be a marking of $\Gamma$. We say that $R$ is also a q-marking of $\Gamma$ if $R_{v}(v)=v$ for all $v\in V$. In this case, we call $(\Gamma,R)$ a q-marked graph.
  • Conversely, we say that a right quasigroup $Q$ is realizable by $\Gamma$ if there exists a marking $R$ of $\Gamma$ such that $Q\cong V_{R}^{\Gamma}$.
Remark 3.4.
Bardakov calls $V_{R}^{\Gamma}$ a graph groupoid. We eschew this name to emphasize the fact that $V_{R}^{\Gamma}$ is not only a magma but also a right quasigroup.
If $V_{R}^{\Gamma}$ is also a rack (resp. quandle), Bardakov calls it a graph rack (resp. graph quandle).
Observation 3.5.
If $(\Gamma,R)$ is a q-marked graph, then $V^{\Gamma}_{R}$ is a rack if and only if $V^{\Gamma}_{R}$ is a quandle.
Observation 3.6.
Since markings of a graph $\Gamma$ are simply functions $R:V(\Gamma)\to\operatorname{Aut}\Gamma$, the number of right quasigroup structures on $V(\Gamma)$ whose right-multiplication maps lie in $\operatorname{Aut}\Gamma$ equals
$\lvert\operatorname{Aut}\Gamma\rvert^{|V(\Gamma)|}.$
Example 3.7.
Given any (di)graph $\Gamma$, the trivial marking $v\mapsto\operatorname{id}_{V(\Gamma)}$ realizes a trivial quandle.
Example 3.8.
Let $n\in\mathbb{Z}^{+}$ be a positive integer, and let $\Gamma$ be the star graph $K_{1,n}$ of order $n+1$; cf. Table 1 in Section 5. Then $\operatorname{Aut}\Gamma\cong S_{n}$ acts on $V:=V(\Gamma)$ by permuting the leaves.
If $n=2$, it is easy to see that of the eight possible markings $R:V\to S_{n}$, exactly four are rack structures. Indeed, if $\ell_{1},\ell_{2}\in V$ are the two leaves and $v\in V$ is the central vertex, then Equation (2.1) forces $R_{\ell_{1}}=R_{\ell_{2}}$. For an example of a marking of $K_{1,2}$ that does not realize a rack, see Example 4.1.
We give more substantial examples following our discussion of Cayley digraphs; see Section 4.
3.2.1. Right quasigroups are realizable
We answer the first question in Problem 1.7. Recall that a digraph $\Gamma=(V,E)$ is called complete if
$E=\{(v,w)\in V\times V\mid v\neq w\}.$
Complete graphs are defined similarly.
Proposition 3.9.
All right quasigroups are realizable by edgeless graphs and complete (di)graphs.
Proof.
Given a right quasigroup $V_{R}$, let $\Gamma$ be an edgeless or complete (di)graph with vertex set $V$. Then $\operatorname{Aut}\Gamma=S_{V}$, so $R:V\to S_{V}$ is a marking of $\Gamma$. Hence, $V_{R}^{\Gamma}=V_{R}$. ∎
Remark 3.10.
In Bardakov’s original wording [bardakov], Proposition 3.9 implies that every groupoid (resp. rack, quandle) is a graph groupoid (resp. graph rack, graph quandle).
3.3. Cayley graphs
Having established Proposition 3.9, it is natural to ask which racks are realized by (di)graphs with more intricate structures. To that end, we discuss Cayley (di)graphs of magmas as introduced by Caucal [caucal].
Definition 3.11.
Let $V_{R}$ be a magma.
  • The (generalized) Cayley digraph of $V_{R}$ with respect to a subset $S\subseteq V$ is the digraph $\Gamma(V_{R},S)$ with vertex set $V$ and edge set
    $E:=\{(v,R_{s}(v))\mid v\in V,s\in S\}\subseteq V\times V.$
    We say that $S$ is the connection set of $\Gamma(V_{R},S)$.
  • The (generalized) Cayley graph of $V_{R}$ with respect to $S$, denoted by $\Gamma_{\mathrm{und}}(V_{R},S)$, is the underlying (simple undirected) graph of $\Gamma(V_{R},S)$.
  • If $S=V$, then we call $\Gamma(V):=\Gamma(V_{R},V)$ the full Cayley digraph of $V_{R}$. The full Cayley graph of $V_{R}$ is defined similarly.
Remark 3.12.
[…]
3.4. Schreier graphs
As Iwamoto et al. [metrics] note, Cayley (di)graphs of right quasigroups are special types of Schreier (di)graphs, which are important objects of study in combinatorial and geometric group theory.
Definition 3.13.
Let $T$ be a subset of a group $G$, and let $V$ be a (left) $G$-set. The (generalized) Schreier digraph $\Gamma^{\mathrm{Sch}}(G,V,T)$ is the digraph with vertex set $V$ and edge set
$E:=\{(v,t\cdot v)\mid v\in V,\,t\in T\}\subseteq V\times V.$
The (undirected) Schreier graph $\Gamma_{\mathrm{und}}^{\mathrm{Sch}}(G,V,T)$ is defined similarly.
Observation 3.14.
For all vertices $v,w\in V$ of $\Gamma:=\Gamma_{\mathrm{und}}^{\mathrm{Sch}}(G,V,T)$, the pair $\{v,w\}$ is an edge of $\Gamma$ if and only if there exists an element $t\in T$ such that $t\cdot w=v$ or $t^{-1}\cdot w=v$.
Remark 3.15.
If $T$ is a symmetric generating subset of $G$, then taking $V:=G$ with the (left) regular action in Definition 3.13 recovers the traditional definition of the Cayley (di)graph of a group.
Remark 3.16.
Given a right quasigroup $V_{R}$ and a subset $S\subseteq V$, let $G:=\operatorname{RMlt}V_{R}$ and T:=R(S){T:=R(S)}. Then
$\Gamma^{\mathrm{Sch}}(G,V,T)=\Gamma(V_{R},S),\quad\quad\Gamma_{\mathrm{und}}^{\mathrm{Sch}}(G,V,T)=\Gamma_{\mathrm{und}}(V_{R},S).$
In the case that $V_{R}$ is a quandle and $R(S)$ generates $G$, Iwamoto et al. called $\Gamma_{\mathrm{und}}(V_{R},S)$ an inner graph; in recent work, they applied the above equality to study quandles using methods from geometric group theory [metrics]*Section 3.
3.4.1. Preliminary results
Our solutions to Problem 1.7 can be stated nicely in terms of Schreier (di)graphs; cf. Theorems 6.16.2.
Proposition 3.17.
Let $\Gamma:=\Gamma^{\mathrm{Sch}}(G,V,T)$ be a Schreier digraph with edge set $E$, and let $H$ be a generating subset of $G$. The following are equivalent:
  1. (1)
    The action of $G$ on $V$ is also an action on $\Gamma$ by digraph automorphisms.
  2. (2)
    For all elements $h\in H$, $v\in V$, and $s\in T$, there exists an element $t\in T$ such that
    (3.1) $th\cdot v=hs\cdot v.$
Proof.
$(1)\implies(2)$: Let $h\in H$, $s\in T$, and $v\in V$, so $(v,s\cdot v)\in E$. By assumption, $(h\cdot v,hs\cdot v)\in E$, so there exists an element $t\in T$ that satisfies Equation (3.1).
$(2)\implies(1)$: To show that $G$ acts on $\Gamma$ by digraph automorphisms, it suffices to show that $(h\cdot v,hs\cdot v)\in E$ for all elements $h\in H$ and directed edges $(v,s\cdot v)\in E$. By assumption, for all such elements and edges, there exists an element $t\in T$ that satisfies Equation (3.1). Hence, $(h\cdot v,hs\cdot v)\in E$. ∎
Proposition 3.18.
Let $\Gamma:=\Gamma_{\mathrm{und}}^{\mathrm{Sch}}(G,V,T)$ be a Schreier graph, and let $H$ be a generating subset of $G$. The following are equivalent:
  1. (1)
    The action of $G$ on $V$ is also an action on $\Gamma$ by graph automorphisms.
  2. (2)
    For all elements $h\in H$, $v\in V$, and $s\in T$, there exists an element $t\in T$ such that one of the following equations holds:
    $th\cdot v=hs\cdot v,\quad\quad h\cdot v=ths\cdot v.$
Proof.
The proof is nearly identical to that of Proposition 3.17; the only difference lies in using Observation 3.14 in the obvious ways. ∎
4. Motivating examples
We consider several examples of Cayley (di)graphs $\Gamma$ of right quasigroups $V_{R}$ and whether or not $(\Gamma,R)$ is a marked graph; see also [bardakov] and [textbook]*Section 1.15. These constructions serve as useful (counter)examples later in the paper.
Example 4.1.
Equip the set $V=[3]$ with the right quasigroup structure given by
$R_{1}=\operatorname{id}_{V},\quad R_{2}=(23),\quad R_{3}=(13).$
By Lemma 2.12, $V_{R}$ is not a rack because
$R_{3}R_{2}R_{3}^{-1}=(12)\notin R(V).$
Figure 1 depicts the full Cayley digraph $\Gamma(V_{R})$ and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
$123$
Figure 1. Full Cayley digraph and full Cayley graph of the right quasigroup from Example 4.1.
Evidently, $V_{R}$ is not realizable by either $\Gamma(V_{R})$ or $\Gamma_{\mathrm{und}}(V_{R},V)$; the only nontrivial (di)graph automorphism is $(12)$. In particular, $R$ is a marking of neither $\Gamma(V_{R})$ nor $\Gamma_{\mathrm{und}}(V_{R},V)$.
Example 4.2.
Equip the set $V=[3]$ with the quandle structure given by
$R_{1}=(23),\quad R_{2}=(13),\quad R_{3}=(12).$
Note that $V_{R}$ is a kei.
Let $S=\{1\}$. Figure 2 depicts the partial Cayley digraph $\Gamma(V_{R},S)$, the full Cayley digraph $\Gamma(V_{R})$, and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
123123123
Figure 2. Partial and full Cayley digraphs and full Cayley graph of the quandle from Example 4.2.
Although $R$ is not a marking of $\Gamma(V_{R},S)$ (or even $\Gamma_{\mathrm{und}}(V_{R},S)$), it is a marking of $\Gamma(V_{R})$ and, hence, of $\Gamma_{\mathrm{und}}(V_{R},V)$.
Example 4.3.
Nonisomorphic right quasigroups may share the same Cayley graphs and even the same Cayley digraphs. For example, let $V$ be the set $[3]$. Figure 3 depicts the full Cayley digraph of the right quasigroup $V_{R}$ defined by
$R_{1}=(12),\quad R_{2}=(13),\quad R_{3}=(23),$
the full Cayley digraph of the permutation rack $V_{(123)}$, and the full Cayley graph shared by $V_{R}$ and $V_{(123)}$.
123123123
Figure 3. Full Cayley digraphs of the two right quasigroups from Example 4.3 and their shared full Cayley graph.
Evidently, $V_{R}$ has the same full Cayley digraph as the quandle from Example 4.2. Moreover, $V_{R}$ and $V_{(123)}$ have the same full Cayley graph. Of course, none of the right quasigroups in question are isomorphic; $V_{R}$ is not a rack, $V_{(123)}$ is a non-involutory rack, and the quandle from Example 4.2 is a kei.
Example 4.4.
Equip the set $V=[4]$ with the right quasigroup structure given by
$R_{1}=\operatorname{id}_{V},\quad R_{2}=(1234),\quad R_{3}=(13)(24),\quad R_{4}=(24).$
Note that
$R_{2}R_{4}R_{2}^{-1}=(13)\notin R(V)\cup R(V)^{-1}.$
In particular, Lemma 2.12 shows that $V_{R}$ is not a rack.
Figure 4 depicts the full Cayley digraph $\Gamma:=\Gamma(V_{R})$ and the full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)$.
$1234$
Figure 4. Full Cayley digraph and full Cayley graph of the right quasigroup from Example 4.4.
Unlike in Example 4.1, $R$ is a marking of $\Gamma$, so $(\Gamma,R)$ realizes $V^{\Gamma}_{R}$. (In fact, not only does $R$ land in $\operatorname{Aut}\Gamma$, but in fact $\operatorname{RMlt}V_{R}=\operatorname{Aut}\Gamma\cong D_{4}$.)
Example 4.5.
Equip the set $V=[5]$ with the quandle structure given by
$R_{1}=(345),\quad R_{2}=(354),\quad R_{3}=(12)(45),\quad R_{4}=(12)(35),\quad R_{5}=(12)(34).$
Let $S=\{1\}$. Figure 5 depicts the partial Cayley digraph $\Gamma=\Gamma(V_{R},S)$ and its underlying graph $\Gamma_{\mathrm{und}}(V_{R},S)$.
$12345$
Figure 5. Partial Cayley digraph and underlying Cayley graph of the right quasigroup from Example 4.5.
Evidently, $R$ is not a marking of $\Gamma(V_{R},S)$, but it is a marking of $\Gamma_{\mathrm{und}}(V_{R},S)$. (Indeed, the automorphism group of the former is $\langle(345),(12)\rangle\cong\mathbb{Z}/6\mathbb{Z}$, which does not contain $R(V)$, while the automorphism group of the latter is $\operatorname{RMlt}V_{R}\cong S_{3}\times\mathbb{Z}/2\mathbb{Z}$.)
5. From marked graphs to racks
In this section, we briefly deduce an answer to Problem 1.6. As an application, we introduce two rack-theoretic invariants of graphs.
5.1. Solution to Problem 1.6
Our alternative characterization of racks does the heavy lifting.
Theorem 5.1.
Let $R$ be a marking (resp. q-marking) of a (di)graph $\Gamma$ with vertex set $V$. Then $V^{\Gamma}_{R}$ is a rack (resp. quandle) if and only if $R$ is a magma homomorphism from $V^{\Gamma}_{R}$ to $\operatorname{Conj}(\operatorname{Aut}\Gamma)$.
Proof.
The statement for markings follows immediately from Proposition 2.11. Thus, the statement about q-markings follows from Observation 3.5. ∎
Remark 5.2.
Theorem 5.1 can be rephrased to state that $(\Gamma,R)$ realizes a rack if and only if the group action of $\operatorname{Aut}\Gamma$ on $V(\Gamma)$ restricts to a rack action of $R(V(\Gamma))$ on $V(\Gamma)$; cf. [survey].
Example 5.3.
Let $\Gamma$ be the complete digraph $\overrightarrow{K_{3}}$ of order 3 (with loops). Then $\Gamma$ is the full Cayley digraph of the quandle from Example 4.2 and the non-rack right quasigroup from Example 4.3.
These two right quasigroups are constructed using the same permutations of $V=[3]$, all of which happen to be automorphisms of $\Gamma$. However, the different choices of vertices that $R:V^{\Gamma}_{R}\to\operatorname{Conj}(\operatorname{Aut}\Gamma)$} assigns to those automorphisms determine whether or not $R$ is a magma homomorphism and, hence, whether or not $V^{\Gamma}_{R}$ is a rack.
5.2. Application of Theorem 5.1
Given a (di)graph $\Gamma$, let $\mu_{\mathrm{rack}}(\Gamma)$ (resp. $\mu_{\mathrm{qnd}}(\Gamma)$) be the number of markings of $\Gamma$ that realize racks (resp. quandles). As an application of Theorem 5.1, we compute these numbers for several graphs and discuss how to compute them in general. This is motivated by Problem 1.6.
In light of Remark 5.2, Theorem 5.1 immediately implies the following.
Corollary 5.4.
Let $\Gamma_{1}$ and $\Gamma_{2}$ be (di)graphs whose automorphism groups are isomorphic to a group $G$. If $V(\Gamma_{1})\cong V(\Gamma_{2})$ as $G$-sets, then $(\mu_{\mathrm{rack}}(\Gamma_{1}),\mu_{\mathrm{qnd}}(\Gamma_{1}))=(\mu_{\mathrm{rack}}(\Gamma_{2}),\mu_{\mathrm{qnd}}(\Gamma_{2}))$. ∎
Corollary 5.5.
The numbers $\mu_{\mathrm{rack}}$ and $\mu_{\mathrm{qnd}}$ are (di)graph invariants. ∎
To compute $\mu_{\mathrm{rack}}(\Gamma)$ given the action of $G:=\operatorname{Aut}\Gamma$ on $V:=V(\Gamma)$, Theorem 5.1 implies that it suffices to count the number of functions $R:V\to G$ that are magma homomorphisms from $V_{R}$ to $\operatorname{Conj}G$. When $n:=|V|$ is finite, this calculation is possible via a computer search. Namely, we use the GRAPE package [GRAPE] in GAP [GAP4] to compute the image $\rho(G)\cong G$ of the permutation representation $\rho:G\hookrightarrow S_{n}$ under the identification $V=[n]$. To compute $\mu_{\mathrm{rack}}(\Gamma)$ and $\mu_{\mathrm{qnd}}(\Gamma)$, we go through all $|G|^{n}$ possible functions $R:V\to\rho(G)$ and count how many of the corresponding right quasigroup structures satisfy the rack and quandle axioms.
We provide an implementation of this exhaustive search algorithm in a GitHub repository [code]. With this implementation, we were able to compute $\mu_{\mathrm{rack}}$ and $\mu_{\mathrm{qnd}}$ for complete graphs $K_{n}$ (equivalently, edgeless graphs), star graphs $K_{1,n-1}$, and cycle graphs $C_{n}$ for small values of $n$; see Table 1. We also have the following general results for path graphs and cycle graphs.
$n$ 0 1 2 3 4 5 6 7
$K_{n}$ $(1,1)$ $(1,1)$ $(2,1)$ $(13,5)$ $(114,36)$ ? ? ?
$K_{1,n-1}$ n/a $(1,1)$ $(2,1)$ $(4,2)$ $(31,13)$ $(390,114)$ ? ?
$C_{n}$ n/a n/a n/a $(13,5)$ $(32,8)$ $(41,7)$ $(108,13)$ $(113,9)$
Table 1. Computations of $(\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}})$ for complete graphs $K_{n}$, star graphs $K_{1,n-1}$, and cycle graphs $C_{n}$ for small values of $n$.
Proposition 5.6.
Let $P_{n}$ be a path graph of order $n\geq 2$. Then
$\mu_{\mathrm{rack}}(P_{n})=\begin{cases}2^{k}&\text{if }n=2k,\\ 2^{k+1}&\text{if }n=2k+1,\end{cases}\qquad\mu_{\mathrm{qnd}}(P_{n})=\begin{cases}1&\text{if }n=2k,\\ 2&\text{if }n=2k+1.\end{cases}$
Proof.
Let $V(P_{n})=[n]$ be the vertices of $P_{n}$ in order, that is, the edges of $P_{n}$ are $(i,i+1)$ for all $1\leq i\leq n-1$. Then the nonidentity element $\sigma$ of $\operatorname{Aut}P_{n}\cong\mathbb{Z}/2\mathbb{Z}$ swaps $i$ and n+1in+1-i for all $1\leq i\leq\lfloor n/2\rfloor$. Therefore, a function $R:V(P_{n})\to\operatorname{Conj}(\operatorname{Aut}P_{n})$ is a magma homomorphism if and only if $R_{i}=R_{n+1-i}$ for all $1\leq i\leq n-1$. Since $\sigma$ has no fixed points if $n=2k$ and a unique fixed point when $n=2k+1$, the claim follows from Theorem 5.1. ∎
Proposition 5.7.
Let $C_{n}$ be a cycle graph of order $n\geq 3$, and let $\sigma(n)$ be the sum of all divisors of $n$. Then
$\mu_{\mathrm{qnd}}(C_{n})=\sigma(n)+1.$
The proof of Proposition 5.7 uses the geometric interpretation of reflections in the dihedral group $D_{n}\cong\operatorname{Aut}C_{n}$}. Although the proof is not terribly long, we defer it to Appendix A to avoid interrupting the flow of the paper.
Example 5.8.
When $n=3$, five markings of the cycle graph $C_{3}$ realize quandles. In particular, let $V_{R}$ be the quandle from Example 4.2. The full Cayley graph $\Gamma_{\mathrm{und}}(V_{R},V)\cong C_{3}$ depicted in Figure 2, marked by the quandle structure $R:V\to S_{3}=\operatorname{Aut}C_{3}$, realizes $V_{R}$. In the following section, we generalize this by showing that all racks are realized by their full Cayley (di)graphs.
6. From racks to marked graphs
In this section, we answer the second question in Problem 1.7. We start by addressing a generalized version of the question for right quasigroups and deduce solutions for racks afterward.
6.1. Results for right quasigroups
Propositions 3.17 and 3.18 do the heavy lifting in the directed and undirected cases, respectively.
Theorem 6.1.
Let $S$ be a subset of a right quasigroup $V_{R}$, and let $\Gamma:=\Gamma(V_{R},S)$. The following are equivalent:
  1. (1)
    $(\Gamma,R)$ is a marked digraph that realizes $V_{R}$.
  2. (2)
    $R$ is a marking of $\Gamma$.
  3. (3)
    For all $h,v\in V$ and $s\in S$, there exists an element $t\in S$ such that
    $R_{t}R_{h}(v)=R_{h}R_{s}(v).$
Proof.
$(1)\iff(2)$: Immediate.
$(2)\iff(3)$: By definition, $H:=R(V)$ generates $G:=\operatorname{RMlt}V_{R}$. In light of Remark 3.16, the equivalence of (2) and (3) is a special case of Proposition 3.17. ∎
Theorem 6.2.
Let $S$ be a subset of a right quasigroup $V_{R}$, and let $\Gamma:=\Gamma_{\mathrm{und}}(V_{R},S)$. The following are equivalent:
  1. (1)
    $(\Gamma,R)$ is a marked graph that realizes $V_{R}$.
  2. (2)
    $R$ is a marking of $\Gamma$.
  3. (3)
    For all $h,v\in V$ and $s\in S$, there exists an element $t\in S$ such that one of the following equations holds:
    $R_{t}R_{h}(v)=R_{h}R_{s}(v),\quad\quad R_{h}(v)=R_{t}R_{h}R_{s}(v).$
Proof.
Similar to the proof of Theorem 6.1, with Proposition 3.18 in place of Proposition 3.17. ∎
Remark 6.3.
The conditions of Theorem 6.2 are strictly weaker than those of Theorem 6.1. Indeed, Example 4.5 gives an example of a quandle $V_{R}$ and a subset $S\subseteq V$ such that $R$ is a marking of $\Gamma_{\mathrm{und}}(V_{R},S)$ but not a marking of $\Gamma(V_{R},S)$.
6.2. Specialization to racks
By considering the third conditions in Theorems 6.16.2, we answer the second question in Problem 1.7 in its original form: All racks are realizable by their full Cayley (di)graphs.
Corollary 6.4.
In the setting of Theorem 6.1 (resp. Theorem 6.2), if conjugating $R(S)$ by elements of $R(V)$ lands in $R(S)$ (resp. $R(S)\cup R(S)^{-1}$), then $(\Gamma,R)$ is a marked digraph (resp. graph) realizing $V_{R}$. In particular, if $V_{R}$ is a rack and $\Gamma$ is its full Cayley (di)graph, then $(\Gamma,R)$ realizes $V_{R}$.
Proof.
The conditions in the first claim directly imply the third conditions in Theorems 6.16.2. Therefore, the second claim follows from Lemma 2.12. ∎
Remark 6.5.
If $V_{R}$ is a right quasigroup but not a rack, then it is not true in general that conjugation in $R(V)$ lands in $R(V)$ (resp. $R(V)\cup R(V)^{-1}$). Nevertheless, the full Cayley digraph (resp. graph) may still satisfy the conditions of Theorem 6.1 (resp. Theorem 6.2); see Example 4.4.
Remark 6.6.
Certainly, the full Cayley (di)graphs of non-rack right quasigroups do not satisfy the conditions of Theorems 6.16.2 in general; see Example 4.1.
Remark 6.7.
The partial Cayley (di)graphs of quandles do not satisfy the conditions of Theorems 6.16.2 in general; see Examples 4.2 and 4.5.
7. Characterization of labeled Cayley digraphs
In this section, we give a graph-theoretic characterization of labeled Cayley digraphs of right-cancellative magmas, right-divisible magmas, right quasigroups, and certain classes of racks.
7.1. Preliminaries
First, we recall several definitions from nonassocative algebra and the theory of labeled digraphs.
7.1.1. Generalizations of right quasigroups
We recall two classes of magmas that generalize right quasigroups.
Definition 7.1.
Let $V_{R}$ be a magma. We say that $V_{R}$ is right-cancellative (resp. right-divisible) if $R_{v}$ is injective (resp. surjective) for all $v\in V$.
Observation 7.2.
A magma is a right quasigroup if and only if it is both right-cancellative and right-divisible.
Observation 7.3.
If $V_{R}$ is a finite magma, then $V_{R}$ is right-cancellative if and only if it is right-divisible.
Example 7.4.
As in Example 2.3, the magma structure given by the addition maps $R_{y}(x):=x+y$ on the positive rational numbers $\mathbb{Q}^{+}$ yields a right-cancellative magma that is not right-divisible.
Example 7.5.
Conversely, the magma structure given by $R_{y}(x):=x^{3}-x$ on the rational numbers $\mathbb{Q}$ yields a right-divisible magma that is not right-cancellative.
7.1.2. Labeled digraphs
[…]
Definition 7.6.
  • A labeled digraph is a triple $\Gamma=(V,E,L)$ where $V$ is a set, $L\subseteq V$ is a subset, and $E\subseteq V\times L\times V$. Given $(v,\ell,w)\in E$, we say that $\ell$ is the label of the edge $(v,\ell,w)$. We say that $V$, $L$, and $E$ are the vertex, labelng, and (labeled) edge sets of $\Gamma$, respectively.
  • The labeled Cayley digraph of a magma $V_{R}$ with respect to a connection set $S\subseteq V$, denoted by $\Gamma_{\mathrm{lab}}(V_{R},S)$, is the labeled digraph $(V,E,S)$ in which
    $E=\{(v,s,R_{s}(v))\mid v\in V,\,s\in S\}.$
Remark $7.$
Labeled edges $(v,\ell,w)\in E$ can also be denoted by transitions $v\xrightarrow{\ell}w$; for example, see Figures 68.
Definition 7.8.
Let $\Gamma=(V,E,L)$ be a labeled digraph.
  • Let $\pi_{1}:E\to V\times L$ and $\pi_{2}:E\to L\times V$ be the projections from $E$ onto its first two and last two coordinates, respectively.
  • We say that $\Gamma$ is deterministic (resp. codeterministic) if $\pi_{1}$ (resp. $\pi_{2}$) is injective.
  • We say that $\Gamma$ is source-complete or executable (resp. target-complete or coexecutable) if $\pi_{1}$ (resp. $\pi_{2}$) is surjective.
Let $\mathcal{D}$ denote the class of deterministic, source-complete labeled digraphs.
The following is immediate.
Observation 7.9.
Let $V_{R}$ be a magma, let $S\subseteq V$, and let $\Gamma=\Gamma_{\mathrm{lab}}(V_{R},S)$. Then:
  • $\Gamma$ lies in $\mathcal{D}$.
  • If $V_{R}$ is right-cancellative, then $\Gamma$ is codeterministic.
  • If $V_{R}$ is right-divisible, then $\Gamma$ is target-complete.
Our first objective will be to prove a converse to Observation 7.9; see Proposition 7.11.
7.2. Construction of $V^{\Gamma}_{R}$
Given an element $\Gamma=(V,E,L)$ of $\mathcal{D}$, we define a magma structure $R$ on $V$ as follows. Since we do not consider marked graphs for the remainder of the paper, we denote this magma by $V^{\Gamma}_{R}$, overwriting the notation from previous sections.
For each non-label vertex $v\in V\setminus L$, let $R_{v}:=\operatorname{id}_{V}$. Otherwise, for each label $\ell\in L$, define $R_{\ell}:V\to V$} as follows. Given a vertex $v\in V$, the preimage $\pi_{1}^{-1}(v,\ell)$ contains a unique element $(v,\ell,w)$ because $\Gamma$ is deterministic and source-complete. So, let $R_{\ell}(v):=w$. The assignment $v\mapsto R_{v}$ makes $V$ into a magma $V^{\Gamma}_{R}$. Verifying the following is straightforward.
Observation 7.10.
If $\Gamma=(V,E,L)$ is an element of $\mathcal{D}$, then $\Gamma$ is the labeled Cayley digraph $\Gamma_{\mathrm{lab}}(V^{\Gamma}_{R},L)$.
7.3. First results
We apply our construction of $V^{\Gamma}_{R}$.
Proposition 7.11.
Let $\Gamma=(V,E,L)$ be an element of $\mathcal{D}$.
  1. (1)
    If $\Gamma$ is codeterministic, then $V^{\Gamma}_{R}$ is right-cancellative.
  2. (2)
    If $\Gamma$ is target-complete, then $V^{\Gamma}_{R}$ is right-divisible.
Proof.
(1): Suppose that $\Gamma$ is codeterministic, so $\pi_{2}$ is injective. We have to show for all $\ell\in V$ that $R_{\ell}$ is injective. If $\ell\notin L$, we are done. Otherwise, suppose for some $v,w\in V$ that
$R_{\ell}(v)=R_{\ell}(w)=:x.$
Since $\pi_{2}$ is injective, $\pi_{2}^{-1}(\ell,x)$ contains at most one element. Since $\pi_{2}^{-1}(\ell,x)$ contains $(v,\ell,x)$ and $(w,\ell,x)$, we obtain $v=w$, as desired.
(2): Suppose that $\Gamma$ is target-complete, so $\pi_{2}$ is surjective. We have to show for all $\ell\in V$ that $R_{\ell}$ is surjective. If $\ell\notin L$, we are done. Otherwise, let $w\in V$. By hypothesis, $\pi_{2}^{-1}(\ell,w)$ contains an edge of $\Gamma$, say $(v,\ell,w)$. Hence, $R_{\ell}(v)=w$. ∎
Henceforth, let $\mathcal{Q}$ denote the subclass of $\mathcal{D}$ whose elements are also codeterministic and target-complete. The following answers Problem 1.8 for right-cancellative magmas, right-divisible magmas, and right quasigroups.
Theorem 7.12.
A labeled digraph is the labeled Cayley digraph of a right-cancellative magma (resp. right-divisible magma, right quasigroup) if and only if it is an element of $\mathcal{D}$ that is codeterministic (resp. target-complete, contained in $\mathcal{Q}$).
Proof.
The first two claims follow from Observations 7.97.10 and Proposition 7.11. Therefore, the third claim follows from Observation 7.2. ∎
7.4. Specialization to racks
Next, we specialize Theorem 7.12 to racks, quandles, involutory right quasigroups, involutory racks, and kei. To that end, we introduce several graph-theoretic conditions corresponding to the labeled Cayley digraphs of objects in these categories.
Recall from Observation 7.10 and Proposition 7.11 that each element $\Gamma=(V,E,L)\in\mathcal{Q}$} is the labeled Cayley digraph of the right quasigroup $V^{\Gamma}_{R}$ with respect to the connection set $L$.
Definition 7.13.
Let $\Gamma=(V,E,L)$ be an element of $\mathcal{Q}$.
  • We say that $\Gamma$ satisfies the first rack condition if for all $v\in V$, if $(v,\ell_{1},w_{1}),(v,\ell_{2},w_{2})\in E$,
    $R_{\ell_{1}}(w_{2})=R_{R_{\ell_{1}}(\ell_{2})}(w_{1}).$
    See Figure 6 for a visualization. Note that we do not assume that $w_{1}\neq w_{2}$.
[…]
    Figure 6. The rack condition from Definition 7.13 states that for all subgraphs of the form on the left, there also exists a subgraph of the form on the right, where $x:=R_{\ell_{1}}(w_{2})$.
  • We say that $\Gamma$ satisfies the second rack condition if for all edges $(v,\ell,w)$ and non-label vertices $x\in V\setminus L$ such that $R_{\ell}(x)\in L$ is a label, then $w$ has a loop labeled by $R_{\ell}(x)$; i.e.,
    $(w,R_{\ell}(x),w)\in E.$
    See Figure 7 for a visualization.
    $R_{\ell}(x)\ellvvwwww$
    Figure 7. The second rack condition from Definition 7.13 states that for all subgraphs of the form on the left and non-label vertices $x\in V\setminus L$, if $R_{\ell}(x)\in L$ is a label, then there exists a subgraph of the form on the right.
  • We say that $\Gamma$ is label-idempotent if $R_{\ell}(\ell)=\ell$ for all $\ell\in L$; that is, each vertex $\ell$ contained in the labeling set $L$ has a loop $(\ell,\ell,\ell)\in E$ labeled by $\ell$.
  • We say that $\Gamma$ satisfies the label-involutory if $R^{2}_{\ell}(v)=v$ for all $v\in V$ and $\ell\in L$; that is, the edge set $E$ can be partitioned into loops and cycles of length 2 having the form
    $\{(v,\ell,w),(w,\ell,v)\}.$
    See Figure 8 for a visualization.
    $\ell\ell\ellvvvvww$
    Figure 8. The label-involutory condition from Definition 7.13 states that subgraphs of the forms on the left and the right partition the edge set.
7.4.1. Result
We answer Problem 1.8 for racks, quandles, and involutory right quasigroups. This also yields answers for involutory racks and kei in the obvious way.
Theorem 7.14.
Let $\Gamma=(V,E,L)$ be a labeled digraph.
  1. (1)
    $\Gamma$ is the labeled Cayley digraph of a rack if and only if $\Gamma$ lies in $\mathcal{Q}$ and satisfies the two rack conditions.
  2. (2)
    $\Gamma$ is the labeled Cayley digraph of a quandle if and only if $\Gamma$ lies in $\mathcal{Q}$, satisfies the two rack conditions, and is label-idempotent.
  3. (3)
    $\Gamma$ is the labeled Cayley digraph of an involutory right quasigroup if and only if $\Gamma$ is a label-involutory element of $\mathcal{Q}$.
Proof.
By Theorem 7.12, we can assume that $\Gamma\in\mathcal{Q}$. As noted before, this inclusion implies that $\Gamma=\Gamma_{\mathrm{lab}}(V^{\Gamma}_{R},L)$.
(1): First, suppose that $V^{\Gamma}_{R}$ is a rack; we show that $\Gamma$ satisfies the rack conditions. For all vertices $v\in V$ and edges $(v,\ell_{i},w_{i})\in E$}, we have $w_{i}=R_{\ell_{i}}(v)$. It follows from Equation (2.1) that
$R_{\ell_{1}}(w_{2})=R_{\ell_{1}}R_{\ell_{2}}(v)=R_{R_{\ell_{1}}(\ell_{2})}R_{\ell_{1}}(v)=R_{R_{\ell_{1}}(\ell_{2})}(w_{1}).$
Hence, $\Gamma$ satisfies the first rack condition. Next, let $(v,\ell,w)\in E$, and let $x\in V\setminus L$ satisfy $R_{\ell}(x)\in L$. Since $R_{x}=\operatorname{id}_{V}$ and $R_{\ell}(v)=w$, Equation (2.1) yields
$R_{R_{\ell}(x)}(w)=R_{R_{\ell}(x)}R_{\ell}(v)=R_{\ell}R_{x}(v)=R_{\ell}(v)=w.$
Since $\Gamma=\Gamma_{\mathrm{lab}}(V^{\Gamma}_{R},L)$, it follows that $(w,R_{\ell}(x),w)\in E$, so $\Gamma$ satisfies the second rack condition.
Conversely, suppose that $\Gamma\in\mathcal{Q}$ satisfies the two rack conditions; we show that $V^{\Gamma}_{R}$ is a rack. We have to verify that
$R_{\ell_{1}}R_{\ell_{2}}(v)=R_{R_{\ell_{1}}(\ell_{2})}R_{\ell_{1}}(v)$
for all $\ell_{1},\ell_{2},v\in V$. If $\ell_{1}\notin L$, we are done. Next, suppose that $\ell_{1}\in L$ and $\ell_{2}\notin L$. If $R_{\ell_{1}}(\ell_{2})\notin L$, we are done. Otherwise, applying the second rack condition to the edge $(v,\ell_{1},R_{\ell_{1}}(v))$ yields the desired equality. Finally, if $\ell_{1},\ell_{2}\in L$, then $(v,\ell_{i},w_{i})\in E$ with $w_{i}:=R_{\ell_{i}}(v)$. Since $\Gamma$ satisfies the first rack property,
$R_{\ell_{1}}R_{\ell_{2}}(v)=R_{\ell_{1}}(w_{2})=R_{R_{\ell_{1}}(\ell_{2})}(w_{1})=R_{R_{\ell_{1}}(\ell_{2})}R_{\ell_{1}}(v).$
(2): By the previous claim, it suffices to show that $V^{\Gamma}_{R}$ is a quandle if and only if $\Gamma$ satifies the label-idempotence condition. But this is clear from the construction of $V^{\Gamma}_{R}$.
(3): Clear from the construction of $V^{\Gamma}_{R}$. ∎
Corollary 7.15.
Let $\mathcal{I}$ be the subclass of $\mathcal{Q}$ whose elements are label-involutory and satisfy the two rack conditions. Then $\mathcal{I}$ (resp. the subclass of $\mathcal{I}$ whose elements are label-idempotent) is precisely the class of labeled Cayley digraphs of involutory racks (resp. kei). ∎
8. Open questions
We conclude by proposing directions for future work.
First, Problem 1.6 motivates the following.
Problem 8.1.
Compute $\mu_{\mathrm{rack}}$ and $\mu_{\mathrm{qnd}}$ for more families of (di)graphs.
Problem 8.2.
Add more entries to Table 1.
A more computationally efficient implementation of the algorithm described in Subsection 5.2 will help in addressing Problems 8.18.2.
The existence of nonisomorphic graphs that satisfy the hypotheses of Corollary 5.4—for example, any non-self-complementary graph $\Gamma$ and its complement $\overline{\Gamma}$—shows that the pair $(\mu_{\mathrm{rack}},\mu_{\mathrm{qnd}})$ is not a complete invariant of graphs. In this light, it is interesting to ask the following.
Problem 8.3.
Under what conditions do nonisomorphic graphs share the same values of $\mu_{\mathrm{rack}}$ and/or $\mu_{\mathrm{qnd}}$ without satisfying the hypotheses of Corollary 5.4?
Finally, Problem 1.8 and recent work applying geometric group theory to quandle theory [metrics] motivate further analogues of Theorem 7.14.
Problem 8.4.
Characterize labeled Cayley digraphs of various classes of racks (e.g., medial racks, Latin quandles, fundamental racks of framed links).
Problem 8.5.
Characterize unlabeled Cayley graphs of right quasigroups, racks, and quandles.
Problem 8.6.
Define and characterize Cayley (di)graphs of classes of racks equipped with extra structure (e.g., generalized Legendrian racks [ta], multi-virtual quandles [virtual], symmetric racks [ta2]).
Appendix A Proof of Proposition 5.7
A.1. Preliminaries
Let $n\geq 3$. Recall that a subgroup of the dihedral group $D_{n}$ is called a reflection subgroup if it is either the trivial subgroup or generated by reflections.
Recall that the automorphism group of the cycle graph $C_{n}$ is isomorphic to $D_{n}$. Therefore, by a 1975 result of Cavior [cavior], proving the following will also prove Proposition 5.7.
Proposition A.1.
Let $\mathcal{S}$ be the set of reflection subgroups of $D_{n}$, and let $\mathcal{M}$ be the set of markings $R$ of $C_{n}$ such that $V^{C_{n}}_{R}$ is a quandle. Then there exists a bijection $\varphi:\mathcal{S}\to\mathcal{M}$.
A.2. Construction of $\varphi$
We construct a function $\varphi:\mathcal{S}\to\mathcal{M}$ geometrically. Given a reflection subgroup $G\in\mathcal{S}$, let $\varphi(G)$ be the marking $R:V\to D_{n}$ defined as follows. For each vertex $v\in V$, if $v$ lies on the axis of a reflection $\psi\in G$ (which is necessarily unique if it exists), then let $R_{v}:=\psi$. Otherwise, let $R_{v}:=\operatorname{id}_{V}$.
Lemma A.2.
If $G\in\mathcal{S}$, then $\varphi(G)\in\mathcal{M}$.
Proof.
By construction, $R_{v}(v)=v$ for all $v\in V$. It remains to show that $R$ is a rack structure on $V$. That is, we have to show that
$R_{v}R_{w}R_{v}^{-1}=R_{R_{v}(w)}$
for all vertices $v,w\in V$. But this is geometrically clear: Since $R_{w}$ is either the identity map or the reflection about the axis $\ell$ containing $w$, the composition $R_{v}R_{w}R_{v}^{-1}$ is either the identity map or the reflection about the axis $R_{v}(\ell)$, i.e., the axis containing $R_{v}(w)$. This transformation is precisely $R_{R_{v}(w)}$. ∎
A.3. Bijectivity of $\varphi$
We construct an inverse map $\varphi^{-1}:\mathcal{M}\to\mathcal{S}$ as follows. Given a quandle structure $R\in\mathcal{M}$, each right-multiplication map $R_{v}$ is either the identity map or a reflection. This is because every rotation in $D_{n}$ has no fixed points. Therefore, defining
$\varphi^{-1}(R):=\operatorname{RMlt}V_{R}=\langle R_{v}\mid v\in V\rangle$
yields a function $\varphi^{-1}:\mathcal{M}\to\mathcal{S}$. Verifying that $\varphi$ and $\varphi^{-1}$ are mutually inverse is straightforward. This completes the proof of Proposition A.1 and, hence, that of Proposition 5.7. ∎